OpenStudy (anonymous):

How many multiples of 6 are between 28 and 280

OpenStudy (anonymous):

multiples of 6 can be written in the form 6k\[28<6k<280\]\[\frac{28}{6}<k<\frac{280}{6}\]

OpenStudy (anonymous):

or Do u knw Ap

OpenStudy (anonymous):

30 , 36 , 42 .......276

OpenStudy (anonymous):

An = a + (n-1)d 276= 30 + (n-1)6 276= 30 + 6n - 6 252 = 6n n = 42

OpenStudy (anonymous):

so..there are 42 multiples....))...hope..this helps..

OpenStudy (unklerhaukus):

\[5\times6=30\] \[50\times6=300\]\[45\times6=270\] \[46\times6=276\] \[46-5=\]

OpenStudy (anonymous):

@UnkleRhaukus u missed one..

OpenStudy (unklerhaukus):

which one did i miss? 28,29,...,277,278,279,280 there is some over hang

OpenStudy (unklerhaukus):

ah i see what i dud wrong, i took away 5 when i should have taken away 4,